You earn $50,000 per year at your job. Suppose the CPI is currently 209, and next year the CPI will be 215. How much money will you need to earn next year to have the same purchasing power as you currently have this year?
Answer #1

The % increase in CPI = (215 - 209)/209 * 100 = 6/209 * 100 = 0.0287 * 100 = 2.87%

Therefore, the income should also increase by 2.87% in order to have the same purchasing power. So, the increase in income required = 2.87% of $50,000 = 2.87/100 * $50,000 = 2.87 * $500 = $1,435.

So, next year I will have to earn $50,000 + $1,435 = $51,435.